# A Survey on State-of-the-art Deep Learning Applications and Challenges

### v5

@@ 4.3.5 Electroencephalography (EEG) Classification
In the work of Abdulwahhab et al. [275], EEG waves’ time-frequency image and raw EEG waves served as input elements for CNN and LSTM models. Two signal processing methods, namely Short-Time Fourier Transform (STFT) and CWT, were employed to generate spectrogram and scalogram images, sized at 77 × 75 and 32 × 32, respectively. The experimental findings demonstrated detection accuracies of 99.57% and 99.26% for CNN inputs using CWT Scalograms on the Bonn University dataset and 99.57% and 97.12% using STFT spectrograms on the CHB-MIT dataset. Similarly, in emotion recognition, several deep learning models have been leveraged with EEG signals. For instance, in [276], a subject-independent emotion recognition model was proposed, which utilizes Variational Mode Decomposition (VMD) for feature extraction and DNN as the classifier. Evaluation against the benchmark DEAP dataset demonstrates superior performance of this approach compared to other techniques in subject-independent emotion recognition from EEG signals. Also, some researchers have also combined EEG signals with facial expression and speech in emotion recognition, as seen in [277], [278], and [279], among others. However, EEG signals can vary significantly across individuals, making it challenging to generalize models across different subjects.
@@ 5 Research Challenges
5.1 Availability and Quality
@@
Building and employing deep learning models face several challenges. The training of deep learning requires a large number of instances (examples) to achieve high accuracy and generalization [280]. Furthermore, the complexity of deep neural networks may lead to overfitting, where the model performs well on training data but fails to generalize on new, unseen data. This phenomenon frequently arises when the models is trained on insufficient data, highlighting the importance of diverse and extensive datasets. However, the data collection and annotation are time consuming and often require domain experts, specialized training and standardization [281]. Moreover, this process is prone to error and has the risk of introducing biases into the dataset which can significantly impact the performance of the trained model. One of the approaches to address this issue is transfer learning. Transfer learning involves the use of a deep learning model (known as pre-trained model) that is trained on a large dataset for solving a specific task (with a small dataset) [282]. The pre-trained model serves as a basis for the model training by fine-tuning the weights of the pre-trained model and adapting it to the new prediction task. This approach helps to mitigate the lack of training data in the target domain. Furthermore, transfer learning reduces computational resources required to train the model and helps faster convergence.
@@
5.2 Interpretability and Explainability
Interpretability and explainability is crucial for building trust and understanding how predictive models make decisions especially in high-stake applications such as healthcare and medical image analysis [286]. However, as deep learning models become more intricate and complex with numerous layers, subnetworks and a large number of parameters, the models are often perceived as a “black box” and difficult to explain in terms of decision-making processes. Therefore, it is crucial for the researchers to focus on methods that provide insights into how a deep learning model performs the prediction and how its decisions are influenced by the input data, making it more transparent and trustworthy. Numerous methods have been proposed for interpreting and explaining the decisions of deep learning models which can be categorized into visualization (feature attribution), model distillation and intrinsic (explainable by itself). Visualization methods involve the use of scientific visualization such as saliency maps or heatmaps to express the explanation by highlighting the degree of association between the inputs and the predictions [287]. The heatmaps identify the saliency of the input features influencing the model’s predictions. The visualization approach is simple and intuitive and can be applied to tabular data and image data. Furthermore, it can be used to identify and debug issues in deep learning models, leading to improved performance and robustness.
5.3 Ethics and Fairness
Deep learning models are increasingly being deployed in making high-stake decision including recruitment [292], criminal justice [293] and credit scoring [294]. There are several advantages of deep learning-based systems in which, unlike humans, machines are able to process vast amounts of data and applications quickly and consistently. However, deep learning-based systems have the risk of being prone to biases present in the data used for training which can lead to unfairness and injustice. Numerous efforts have been made to mitigate this issue which can be categorized into modelling bias detection and modelling bias mitigation. Detection of modelling bias refers to the process of identifying and quantifying biases that may present in predictive models. This approach involves the use of statistical analysis, fairness metrics, counterfactual testing and human review to detect bias in the models. For instance, visualization-based methods such as attribution maps are used to indicate which regions are significant to the predictions [295]. This in turn can be used to detect and quantify bias using metrics such as Relevance Mass Accuracy, Relevance Rank Accuracy Accuracy and or Area over the perturbation curve (AOPC). In [296], two modules are presented for estimating bias in predictive models. The first module utilizes an unsupervised deep neural network with a custom loss function to generate hidden representation of the input data called bias vectors, revealing the underlying bias of each feature. The second module combines these bias vectors into a single vector representing the bias estimation of each feature, achieved by aggregating them using the absolute averaging operation. Bias mitigation refers to the process of reducing the presence of bias in predictive models, which can be done in three stages. The first stage combats bias by modifying the training data, either relabeling the labels or perturbing the feature values [297, 298]. The second stage addresses bias during the training of the model by applying regularization terms to the loss function to penalize discrimination. In [299], a loss function based on bias parity score (BPS) is introduced to measure the degree of similarity of a statistical measure such as accuracy across different subgroups. The BPS term is added to the loss function as a regularizer to the original prediction task. The last stage mitigates bias after the predictive models have been successfully trained. This stage applies post-processing approaches such as reinforcement learning to obtain a fairer model [300]. For instance, the detection of minority classes is rewarded to prevent bias towards the majority class. This allows the model the generalize well across different patient demographics.
5.4 Lightweight Deep Learning Models
Even though deep learning architectures have achieved state-of-the-art across various computer vision tasks, they often come with large model parameters [197]. The architecture and complexity of a deep learning network determine the number of model parameters. The deeper the network, the larger the number of model parameters. However, deep learning models with large parameters often suffer limitations when deploying on end devices. For instance, a deep learning model developed for security monitoring by analyzing video data using 3D-CNNs might suffer deployment issues when deploying such models on low-resourced systems like smartphones or small-scale IoT devices. Model training and inference for deep learning models with large parameters demands substantial processing power. As the number of parameters increases, so does the computational complexity, resulting in longer model training duration and more hardware needs. Also, large parameter sizes translate to increased memory requirements, limiting their deployment on end devices. This is because these end devices often have battery, processor or memory capacity limitations. To address these challenges, it is important to develop sophisticated but lightweight architectures that can achieve state-of-the-art with few model parameters. Such lightweight models will be characterized by their ability to deliver competitive performance while mitigating computational complexity and memory requirements, making them well-suited for deployment on resource-constrained devices. An approach would be to develop novel lightweight plug-and-play modules that can be plugged to few layered deep learning architectures to improve feature learning without incurring additional model complexity. Other approaches could involve leveraging model compression techniques to reduce the size and computational complexity of deep learning models. Researchers can focus on improving pruning methods [301], which can identify and eliminate redundant parameters or connections, thereby reducing the model’s footprint without compromising performance. Also, quantization techniques [302] can be further explored to reduce the precision of weights and activations, therefore, enabling efficient representation with lower memory requirements. Also, knowledge distillation techniques [303] can be further investigated to facilitate the transfer of knowledge from a complex teacher model to a simpler student model, therefore, enabling compact yet effective representations. These areas are still open to contributions.
5.5 Adversarial Attack and Defense
Adversarial attacks and defense mechanisms in deep learning represent a critical area of research and development, particularly as deep learning models become increasingly integrated into various applications. Adversarial attacks involves the deliberate manipulation of input data to mislead or deceive deep learning models, leading to incorrect predictions or behavior [304]. Szegedy et al. [305] was the first to identify this intriguing shortcoming of deep neural networks in image classification. They showed that even with their great accuracy, deep learning models are surprisingly vulnerable to adversarial attacks that take the form of tiny image changes that are (almost) invisible to human vision systems. A neural network classifier may radically alter its prediction about an image as a result of such an attack. Also, such a model can indicate high confidence in wrong predictions, which can be catastrophic for deep learning models deployed in medical or security fields, among many others. In generative models, several studies have investigated how adversarial attacks affect autoencoders and GANs, as seen in Tabacof et al. [306] where a method to manipulate input images in a way that deceives variational autoencoders into reconstructing a totally different image was introduced. In recent times, the focus of adversarial attack research has been on images, but studies has shown that adversarial attacks are not limited to image data; they can also affect other types of data such as text, signals, audio, and video [307, 308, 309].

### v6

@@ 4.3.5 Electroencephalography (EEG) Classification
Hybrid Models: In [331], the authors developed three deep learning architectures (CNN, LSTM, and hybrid CNN-LSTM), with each model chosen for its effectiveness in handling the intricate characteristics of EEG data for epilepsy detection. Each architecture offers distinct advantages, with CNN excelling in spatial feature extraction, LSTM in capturing temporal dynamics, and the hybrid model combining these strengths. The CNN model, consisting of 31 layers, attained the highest accuracy, achieving 91% on the first benchmark dataset and 82% on the second dataset using a 30-second threshold, selected for its clinical significance. In the work of Abdulwahhab et al. [332], EEG waves’ time-frequency image and raw EEG waves served as input elements for CNN and LSTM models. Two signal processing methods, namely Short-Time Fourier Transform (STFT) and CWT, were employed to generate spectrogram and scalogram images, sized at 77 × 75 and 32 × 32, respectively. The experimental findings demonstrated detection accuracies of 99.57% and 99.26% for CNN inputs using CWT Scalograms on the Bonn University dataset and 99.57% and 97.12% using STFT spectrograms on the CHB-MIT dataset. Similarly, in emotion recognition, several deep learning models have been leveraged with EEG signals. For instance, in [333], a subject-independent emotion recognition model was proposed, which utilizes Variational Mode Decomposition (VMD) for feature extraction and DNN as the classifier. Evaluation against the benchmark DEAP dataset demonstrates superior performance of this approach compared to other techniques in subject-independent emotion recognition from EEG signals.
Some researchers have also combined EEG signals with facial expression and speech in emotion recognition, as seen in [334], [335], and [336], among others. In [335], a multi-modal emotion recognition framework is proposed, utilizing three deep learning models to extract features from facial expressions, speech and EEG signals. To process EEG signals, the authors designed a tree-like LSTM model that extracts temporal features at different stages to capture multiscale feature representations. The output of each model is then weighted to generate the final predictions. In [336], the feature maps extracted by the deep learning models are combined through concatenation, and then they are fed into a fully connected classifier to generate the predictions. For EEG signals, two deep learning models are designed: one for extracting local features and the other for extracting global features. Both models are implemented using convolutional layers. However, EEG signals can vary significantly across individuals, making it challenging to generalize models across different subjects.
Generative Models: GANs have been employed to enhance EEG signal classification. In [337], a conditional GAN model is proposed to overcome the limited number of EEG samples. To improve data generation quality, the discriminator is integrated with SE attention and a cropped training strategy is employed to leverage the entire spectrum of the EEG signals. In [338], a Wasserstein GAN model is employed to enhance the spatial resolution of EEG signals for EEG classification. The model takes low-resolution signals and generates the corresponding high-resolution signals while simulataneously interpolating the missing channels. EEG signals are often corrupted with noise and artifacts. DHCT-GAN is a dual-branch hybrid model designed to generate denoised EEG signals by processing both clean and noisy/artifact-corrupted signals [339]. The generator consists of dual-branch feature learning pipelines, where one accepts clean signals while the other handles noisy/artifact signals to learn feature representations. The feature maps are then fused using two fully connected layers with a tanh activation function to generate denoised EEG signals. Both feature learning branches are identical, consisting of convolutional layers with local and global self-attention mechanisms, while the discriminators are implemented using convolutional and fully connected layers. Table 13 lists the summary of state-of-the-art EEG classification.
Table 13: Summary of state-of-the-art EEG classification.
Ref. Description Results (Datasets and Metrics)
[331], 2024 A hybrid CNN-LSTM model consists of a convolutional layer to extract local features and two LSTM layers for temporal modeling UPenn and Mayo Clinic’s Seizure Detection Challenge
Accuracy: 89%
F1-score (avg.): 87%
[335], 2024 A tree-like LSTM to capture multiscale temporal feature representations from EEG signals MAHNOB-HCI
Accuracy: 0.9450 (valence), 0.7628 (arousal)
[336], 2023 Parallel convolutional pipelines for extracting local and global features from EEG signals DEAP
Accuracy: 85.72% (valence), 87.97% (arousal)
MAHNO B-HCI
Accuracy: 85.98% (valence), 85.23% (arousal)
[332], 2024 A framework comprises two separated deep learning-based models using a 2D Convolutional Neural Network and an LSTM to process EEG signal images and normalized EEG signals Bonn
Accuracy: 99.75%
CHB-MIT
Accuracy: 97.12%
[337], 2024 A conditional GAN integrated with SE attention in the discriminator and employing a cropped training strategy for improved data generation BCI Competition IV 2a
Accuracy: 81.3%
BCI Competition IV 2b
Accuracy: 90.3%
[338], 2025 A Wasserstein GAN model is used to enhance the spatial resolution of EEG signals for classification by generating high-resolution signals from low-resolution inputs while simultaneously interpolating missing channels, addressing noise and artifact contamination Dataset V (Berlin BCI)
Accuracy: 83.88% (scale 2), 82.00% (scale 4)
[339], 2025 A dual-branch hybrid model that denoises EEG signals by separately learning features from clean and noisy inputs, fusing them through fully connected layers, and leveraging convolutional layers with self-attention mechanisms EEGdenoiseNet(EEG, EOG, EMG), MIT-BIH Arrhythmia (ECG) and semi-simulated EEG/EOG
$\eta$ (artifact reduction): 82.35% (EMG), 91.80% (EOG), 88.97% (ECG), 82.07% (EOG+EMG)
4.4 Robotics
Deep learning is a crucial component in modern robotics, enabling machines to perceive and interpret their surroundings. By leveraging sensors and cameras, robots can identify objects, navigate environments and interact with humans more effectively.
4.4.1 Object Identification
Deep learning plays a crucial role in object identification, allowing robots to recognize, classify, and differentiate between various objects in real-time. In [340], an intelligent system, integrating deep learning with robotic kinematic control is proposed for waste classification. The system consists of a robotic arm with a gripper and a camera placed in front of the robot to capture real-time video. YOLO model is trained to recognize and localize objects, allowing the system to determine the location, pick them up and classify them. A similar system is reported in [341], where a robotic arm is used to pick up objects on a conveyor and a camera is mounted above the conveyor to capture real-time video. The waste sorting process is divided into two tasks: localization of the waste objects and their classification. The object localization is performed using Segment Anything architecture (SAM) [342] which is based on vision transformer. Several lightweight SAM models are proposed to solve the localization task efficiently while minimizing computational requirements. For classification, object images are cropped using SAM’s output masks and classified by a MobileNetV2.
A deep learning-based method is proposed to detect tile edges and corners for tile-paving mobile robotic systems [343]. The camera is mounted on a tile-grabing platform that is parallel to the ground and approximately 10 to 20 cm above it. The system utilizes a YOLOv8 for semantic segmentation with rule-based post-processing techniques. An unmanned ground vehicle-based is proposed for automated structural damage inspection [344]. The robotic system utilizes LiDAR and a camera, integrating them via Robot Operating System for efficient control and data processing. The images are processed for the damage detection and segmentation using an improved YOLOv7 model, while the point cloud data is processed for real-time localization and 3D mapping using an enhanced Keep It Small and Simple-Iterative Closest Point algorithm. A similar system is reported in [345], where object detection and tracking approach is proposed to fuse point cloud and visual data. The approach leverage YOLOv8 real-time object detection capabilities, while LiDAR provides accurate 3D spatial information and distance measurements, enhancing perception reliability. A fusion method aligns LiDAR data with camera images, enabling accurate 3D bounding boxes and object tracking. Table 14 lists the summary of state-of-the-art robot object identification.
Table 14: Summary of state-of-the-art robot object identification.
Ref. Hardware platform Model Task
[340], 2025 Robotic arm, AI Stereo ZED 2 camera, gripper, Intel Xeon E5 series CPU @ 2.40 GHz YOLOv4 Waste classification
[341], 2025 PC workstation, industrial camera, AMD Threadripper 3970X 32 cores CPU @ 3.79 GHz, 128 GB RAM, two Titan RTX 24 GB GPUs Segment Anything architecture (SAM) Waste sorting
[342], 2023 Mobile robot, camera, Nvidia RTX 3060 GPU YOLOv8 Tile edges and corners detection
[344], 2025 Husky Unmanned ground vehicle, Intel Realsense camera, Velodyne VLP-16 LiDAR, Intel i7 CPU YOLOv7 Structural damage inspection
[345], 2024 Agilex Scout Mini robot, Velodyne LiDAR, Intel D435 camera YOLOv8 Object tracking
4.4.2 Path Extraction and Navigation
Deep learning models have been applied for robot navigation systems. In [346], LiDAR and a camera are to capture point cloud and visual data for a comprehensive understanding and navigation in the environments. The study experiments with Faster R-CNN, YOLOv5 and YOLOv8 for object detection. Additionally, the paper compares the performance of these systems in various real-world environments, showcasing their potential to enhance autonomous navigation. Mobile robot navigation system is proposed utilizing semantic segmentation to determine the drivable paths [347]. First, Deeplabv3 [348] model with a ResNet-50 backbone is applied to segment scene images to extract drivable area. A perspective transformation then maps the segmented images to real-world space. Following the transformation, the image is divided into grids to determine the optimal path while avoiding obstacles. Finally, a PID controller guides the robot along the smoothed path to ensure accurate navigation in the environment.
In [349], a method for orchard robot navigation utilizing a modified YOLOv8 model is proposed. The model locates vine tree trunks and identifies the center points of tree trunks at the lower end of the detection boxes. The least squares method is then applied to fit reference lines on both sides of the trunk. Furthermore, a multiscale attention mechanism is incorporated into the model to prioritize relevant features, enhancing the model performance. In [350], deep learning is employed for navigation path extraction for agricultural robots. The authors propose Single-Stage Navigation Path Extraction Network (NPENet) to predict the road centerline in a single stage, eliminating the need for complex multi-stage processes such as line detection and segmentation. The proposed model uses residual blocks for feature extraction which includes batch normalization, leaky ReLU and convolutional layers, generating three key outputs: detecting if the robot is on the road, estimating the navigation line’s angle and predicting the vanishing point of the road. These outputs enable the robot to determine its trajectory and real-time navigation in orchard environments. Table 15 lists the summary of path extraction and navigation.
Table 15: Summary of state-of-the-art path extraction and navigation.
Ref. Hardware platform Model Task
[346], 2024 Mobile robot, 2MP Logitech webcam, LiDAR, Raspberry Pi 4, 4 GB RAM YOLOv8 Semantic navigation
[347], 2024 Mobile robot, camera, Nvidia Jetson Nano Deeplabv3+ (ResNet50 backbone) Path planning
[349], 2024 Mobile robot, Intel RealSense D455 camera, Intel Core i7 10870H CPU @ 2.20 GHz, Nvidia GeForce GTX 1650 Ti GPU YOLOv8 Navigation line extraction
[350], 2025 Mobile robot, Intel RealSense cameras, Nvidia Jetson Xavier NPENet Navigation path extraction
4.4.3 Human-Robot Interaction
Deep learning has significantly advanced human-robot interaction, particularly through the integration of large language models (LLMs) designed for natural language processing. Despite their extensive knowledge, LLMs have limitations, as they can sometimes generate inaccurate information, a phenomenon known as hallucination [351]. To address this issue, LLMs need to be connected to the physical world by providing real-world contextual information to the models, enabling them to generate accurate responses that are relevant to the situation. This process is known as LLM grounding. In [352], the authors explore the use of LLMs in human-robot interaction, focusing on collaborative planning between humans and robots. The LLM is guided to generate outputs within the predefined set of robot skills using a few-shot prompting, while vision models such as SAM-CLiP and ViLD are used for object identification within the scene, enabling the robot to perform actions in the real-world. By leveraging LLMs, robots can understand and process natural language instructions and execute the corresponding actions, enabling efficient collaboration with human users. LLMs need to be aligned with human intentions to ensure effective communication and interaction between humans and robots. In [353], a 3D scene graph is employed to represent objects and environments captured by 3D instance segmentation and 2D image classification. Given the 3D scene graph, a complex task is decomposed using LLM through step-by-step reasoning by progressively retrieving relevant nodes from the graph. During the task solving, human intent is automatically identified, and the system generates responses accordingly.
In [354], a modular system architecture that can be easily extended is proposed to integrate LLM models for human-robot conversation and collaboration. Figure 33 shows the system architecture, where the chat manager is the central to the system that coordinates the state and inference of the LLM model. The robot perception such as recognizing objects and human gesture is handled separately by different deep learning models while robot actions are operated by speech generation model and actuators for emotion expression, gaze control and arm movement. This approach allows the robot to engage in open-ended conversation and collaborate naturally with humans. Grounding LLMs can be challenging due to robot’s multi-modal perception system. Furthermore, the incoming data has different sampling rate which makes data alignment difficult, causing the robotic system to miss valuable information. In [355], a framework is proposed for processing multi-modal inputs to generate coherent narratives about robot experiences including its internal status, current observation, planning and critical alert. The framework is divided into three stages: key event selection, experience summarization and narration generation modules. First, multi-modal data across three categories: environment, internal and planning are aligned by sampling at a fixed rate and associating each frame at the nearest timestamp. Based on the aligned data, key events are selected by detecting changes in optical flow, RGB images and joint states using heuristic and normalization techniques to identify moment of interest. Using the detected events, the robotic data is converted into natural language summaries, and an LLM generates user-friendly explanation by filtering repetitive and irrelevant information. Table 16 lists the summary of human-robot interaction.
Figure 33: A grounded chat architecture [354].
Table 16: Summary of state-of-the-art human-robot interaction.
Ref. Hardware platform Model Task
[352], 2025 Niryo Ned robot and Intel RealSense D435i camera LLM: GPT-3.5,
Object detection: SAM-CLiP and ViLD
Robot planning
[353], 2024 N.A. LLM: GPT-4 turbo Human collaboration
[354], 2024 Neuro-Inspired COLlaborator (NICOL) robot LLM: GPT, Vicuna, Mistral
Human detection: YOLOX
Pose estimation: HRNet
Object detection: ViLD
Conversation and collaboration
[355], 2024 MStretch SE3 robot, Intel RealSense D435i and D405 Object segmentation: YOLO World Robot behavior announcement
@@ 5 Research Challenges
Despite the success of deep learning in various applications, there are still fundamental challenges faced by researchers and practitioners. This section discusses the challenges ethical, technical and domain-specific perspectives.
5.1 Ethical Issues
Interpretability and explainability is crucial for building trust and understanding how predictive models make decisions, especially in high-stake applications such as healthcare and medical image analysis [356]. However, as deep learning models become more intricate and complex with numerous layers, subnetworks and a large number of parameters, the models are often perceived as a “black box” and difficult to explain in terms of decision-making processes. Therefore, it is crucial for the researchers to focus on methods that provide insights into how a deep learning model performs the prediction and how its decisions are influenced by the input data, making it more transparent and trustworthy. Numerous methods have been proposed for interpreting and explaining the decisions of deep learning models, which can be categorized into visualization (feature attribution), model distillation and intrinsic (explainable by itself). Visualization methods involve the use of scientific visualization such as saliency maps or heatmaps to express the explanation by highlighting the degree of association between the inputs and the predictions [357]. The heatmaps identify the saliency of the input features influencing the model’s predictions. The visualization approach is simple and intuitive and can be applied to tabular data and image data. Furthermore, it can be used to identify and debug issues in deep learning models, leading to improved performance and robustness.
Deep learning models are increasingly being deployed in making high-stake decision including recruitment [362], criminal justice [363] and credit scoring [364]. There are several advantages of deep learning-based systems in which, unlike humans, machines are able to process vast amounts of data and applications quickly and consistently. However, deep learning-based systems have the risk of being prone to biases present in the data used for training which can lead to unfairness and injustice. Numerous efforts have been made to mitigate this issue which can be categorized into modelling bias detection and modelling bias mitigation. Detection of modelling bias refers to the process of identifying and quantifying biases that may present in predictive models. This approach involves the use of statistical analysis, fairness metrics, counterfactual testing and human review to detect bias in the models. For instance, visualization-based methods such as attribution maps are used to indicate which regions are significant to the predictions [365]. This in turn can be used to detect and quantify bias using metrics such as Relevance Mass Accuracy, Relevance Rank Accuracy and Area over the perturbation curve (AOPC). In [366], two modules are presented for estimating bias in predictive models. The first module utilizes an unsupervised deep neural network with a custom loss function to generate hidden representation of the input data called bias vectors, revealing the underlying bias of each feature. The second module combines these bias vectors into a single vector representing the bias estimation of each feature, achieved by aggregating them using the absolute averaging operation.
Bias mitigation refers to the process of reducing the presence of bias in predictive models, which can be done in three stages. The first stage combats bias by modifying the training data, either relabeling the labels or perturbing the feature values [367, 368]. The second stage addresses bias during the training of the model by applying regularization terms to the loss function to penalize discrimination. In [369], a loss function based on bias parity score (BPS) is introduced to measure the degree of similarity of a statistical measure such as accuracy across different subgroups. The BPS term is added to the loss function as a regularizer to the original prediction task. The last stage mitigates bias after the predictive models have been successfully trained. This stage applies post-processing approaches such as reinforcement learning to obtain a fairer model [370]. For instance, the detection of minority classes is rewarded to prevent bias towards the majority class. This allows the model to generalize well across different patient demographics.
5.2 Technical Issues
Even though deep learning architectures have achieved state-of-the-art across various computer vision tasks, they often come with large model parameters [371]. The architecture and complexity of a deep learning network determine the number of model parameters. The deeper the network, the larger the number of model parameters. However, deep learning models with large parameters often suffer limitations when deploying on end devices. For instance, a deep learning model developed for security monitoring by analyzing video data using 3D-CNNs might suffer deployment issues when deploying such models on low-resourced systems like smartphones or small-scale IoT devices. Model training and inference for deep learning models with large parameters demands substantial processing power. As the number of parameters increases, so does the computational complexity, resulting in longer model training duration and more hardware needs. Furthermore, large parameter sizes translate to increased memory requirements, limiting their deployment on end devices. This is because these end devices often have battery, processor, or memory capacity limitations. To address these challenges, it is important to develop sophisticated but lightweight architectures that can achieve state-of-the-art with few model parameters. Such lightweight models will be characterized by their ability to deliver competitive performance while mitigating computational complexity and memory requirements, making them well-suited for deployment on resource-constrained devices. An approach would be to develop novel lightweight plug-and-play modules that can be plugged to a few layered deep learning architectures to improve feature learning without incurring additional model complexity. Other approaches could involve leveraging model compression techniques to reduce the size and computational complexity of deep learning models. Researchers can focus on improving pruning methods [372], which can identify and eliminate redundant parameters or connections, thereby reducing the model’s footprint without compromising performance. Furthermore, quantization techniques [373] can be further explored to reduce the precision of weights and activations, therefore, enabling efficient representation with lower memory requirements. Also, knowledge distillation techniques [374] can be further investigated to facilitate the transfer of knowledge from a complex teacher model to a simpler student model, therefore, enabling compact yet effective representations.
Deep learning training involves complex processes that require efficient optimization to ensure fast convergence and resource management. As models grow in size and complexity, challenges such as slow convergence, vanishing gradients, and computational limitations become more pronounced. Researchers have explored various methods to overcome these issues, proposing innovative approaches like predicting parameter change and incorporating them into training, thus reducing training time and improving the model performance [375]. Researchers also leverage prior experience to optimize parameter adjustments to reduce forward and back propagation steps, thus reducing computational costs and making training more efficient [376]. Others have proposed alternative training algorithms such as alternating minimization [377] and random search based on the annealing method [378]. Overall, the researchers aim to accelerate learning and improve efficiency either by decomposing the optimization problem into sequential sub-problems or by refining the search space based on parameter value ranges.
Adversarial attacks and defense mechanisms in deep learning represent a critical area of research and development, particularly as deep learning models become increasingly integrated into various applications. Adversarial attacks involve the deliberate manipulation of input data to mislead or deceive deep learning models, leading to incorrect predictions or behavior [379]. Szegedy et al. [380] was the first to identify this intriguing shortcoming of deep neural networks in image classification. They showed that even with their great accuracy, deep learning models are surprisingly vulnerable to adversarial attacks that take the form of tiny image changes that are (almost) invisible to human vision systems. A neural network classifier may radically alter its prediction about an image as a result of such an attack. Furthermore, such a model can indicate high confidence in wrong predictions, which can be catastrophic for deep learning models deployed in medical or security fields, among many others. In generative models, several studies have investigated how adversarial attacks affect autoencoders and GANs, as seen in Tabacof et al. [381] where a method to manipulate input images in a way that deceives variational autoencoders into reconstructing an entirely different image was introduced. In recent times, the focus of adversarial attack research has been on images, but studies have shown that adversarial attacks are not limited to image data; they can also affect other types of data such as text, signals, audio, and video [382, 383, 384].
5.3 Domain-specific Issues
@@
Building and employing deep learning models face several challenges. The training of deep learning requires a large number of instances (examples) to achieve high accuracy and generalization [385]. Furthermore, the complexity of deep neural networks may lead to overfitting, where the model performs well on training data but fails to generalize on new, unseen data. This phenomenon frequently arises when the models are trained on insufficient data, highlighting the importance of diverse and extensive datasets. However, the data collection is time-consuming and often require domain experts, specialized training and standardization [386]. Moreover, this process is prone to error and has the risk of introducing biases into the dataset which can significantly impact the performance of the trained model. One of the approaches to address this issue is transfer learning. Transfer learning involves the use of a deep learning model (known as pre-trained model) that is trained on a large dataset for solving a specific task (with a small dataset) [387]. The pre-trained model serves as a basis for the model training by fine-tuning the weights of the pre-trained model and adapting it to the new prediction task. This approach helps to mitigate the lack of training data in the target domain. Furthermore, transfer learning reduces computational resources required to train the model and helps faster convergence.
@@
Data annotation is expensive and time-consuming. The problem is exacerbated when the data is of low quality corrupted with noise which may lead to bad, resulting in unreliable training data. One of the approaches addresses this issue is active learning. Active learning is an approach where the learning algorithm selectively queries the most informative data points for labeling. The aim is to improve predictive model efficiency with fewer training data, thus reducing the overall cost of data annotation. The approach involves training a model using an initial training set, and then a subset of unlabeled data is selected for labeling by external annotators. The newly labeled instances are appended to the training set for retraining or fine-tuning. This process is repeated until the model performance reaches a desired threshold or the labeling budget has exhausted. The core principle behind active learning is the query strategy or, essentially how to select the most informative data for labeling that will be beneficial to the model training. In general, the query may select the most ambiguous instances based on the model predictions, instances that will have a significant impact on the model performance, or instances that cover the distribution of the entire feature space [391]. The dynamic and sequential nature of real-world applications presents different challenges, which led to the development of online active learning. This approach was introduced to address specific issues, including data streams, concept drift and environmental changes [392]. Recently, researchers have investigated the use of LLMs for data annotation. The study presents methods for generating annotations, assessing annotations and utilizing annotations [393].
Self-supervised learning is another approach that was introduced to mitigate the issue of limited labeled datasets. In self-supervised learning, the model learns from unlabeled data by generating its own labels from the input data [394]. This approach essentially creates a pretext task which the model solves without requiring manual annotations or labeled datasets. By solving the pretext task, the model leverages the underlying structure in the data and learn useful representations that can be used for specific tasks. For example, in computer vision, a model might predict the speed [395] or repeated scenes [396] in a video. In natural language processing, it could predict missing words in sentences to learn the dependencies between words [397]. The key advantage of self-supervised learning is that it allows the model to take advantage of large amounts of unlabeled data, which is often much more abundant and cheaper to collect than labeled data. The learned representations can then be used for downstream tasks like classification, regression, or segmentation.